Circular No. 5978

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

FG SAGITTAE E. F. Guinan and G. P. McCook, Villanova University, report: “FG

Sge, the peculiar, pulsating central star of the planetary nebula He 1-5, is undergoing another large dimming event. UBVRI photoelectric photometry, carried out with the Four College Consortium 0.8-m APT at Mt. Hopkins, shows a ∼1.2-mag decrease in its V -band brightness from Mar. 23 to Apr. 19 UT. FG Sge had undergone its first recorded sudden dimming and reddening episode in Aug-Sept. 1992 (IAUC 5604, 5609, 5610, 5619, 5632) and since then has been slowly recovering in brightness. The behavior of FG Sge is probably due to the formation of dust in its outer envelope. Its characteristics since 1992 are similar to those of R CrB-type variables.”

NOVA CASSIOPEIAE 1993 Further visual magnitude estimates (cf. IAUC 5954): Mar. 30.78 UT,

15.9 (L. Szentaskó, Veresegyház, Hungary); Apr. 4.80, 15.9 (Szentaskó); 14.853, [14.4 (T. Vanmunster, Landen, Belgium); 15.09, 14.2: (Szentaskó); 19.09, 13.7 (Szentaskó); 19.115, 13.9 (Vanmunster); 21.07, 10.8 (Szentaskó).

SUPERNOVA 1994I IN NGC 5194 C. J. Chandler, National Radio Astronomy Observatory (NRAO); J. A.

Phillips, California Institute of Technology; and M. P. Rupen, NRAO, report: “Using the five-element millimeter interferometer at the Owens Valley Radio Observatory, we observed SN 1994I at 99 GHz on Apr. 16.24– 16.56 UT. No source was detected, the rms noise in the image being 1.0 mJy/beam. The absolute flux scale was established by observations of Neptune and Uranus; the systematic uncertainty is ∼ 15 percent. The supernova flux density at 3 mm has declined by at least a factor of 3.7 over 11 days (cf. IAUC 5968); assuming an explosion date of Mar. 31 and a power-law decline, this implies a decay proportional to t^{−1} or steeper. This is consistent with the rapid radio decline seen in type-Ib/c supernovae (Weiler et al. 1986, Ap.J. 301, 790; Van Dyk et al. 1993, Ap.J. 419, L69).”

3C 279 A. W. Neely, Silver City, NM; and M. Harvanek and J. Stocke, Univer-

sity of Colorado, report that this quasar is brightening; recent photometry: Apr. 19 UT, V = 14.6 ± 0.2; 21, 14.75 ± 0.15; 22, 14.57 ± 0.05.

1994 April 23 Daniel W. E. Green